OpenStudy (anonymous):

solve the following trigonometric equation: 3sin^(2)x=cos^(2); where x is greater than or equal to 0 and less than 2 pi

myininaya (myininaya):

Put everything on one side and write in terms of one trig function

myininaya (myininaya):

You will need the identity \[\cos^2(x)=1-\sin^2(x)\]

OpenStudy (anonymous):

is it asking for an actual number?

myininaya (myininaya):

Yes we are looking for values of x that satisfy that equation I think you are missing an x \[3 \sin^2(x)=\cos^2(x) \] right?

OpenStudy (anonymous):

Oops! Yes, I was missing an x. So you place everything on one side and set it equal to 0?

myininaya (myininaya):

yes

myininaya (myininaya):

Also write in terms of one trig function

OpenStudy (anonymous):

Wait, would you set it equal to zero or one? Because if it is equal to zero then how do I use the identities?

myininaya (myininaya):

I said put everything on one side. How do you do that? Subtract cos^2(x) on both sides And then use the identity I gave you above Write what you get after you have follow these instructions No about the equals to 1. That wouldn't help (adding a 1 on both sides would not help)

OpenStudy (anonymous):

I got the answer. Thanks